Circular No. 8097

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2003cg, 2003ch, AND 2003ci Three new apparent supernovae have been discovered. SN 2003cg in

NGC 3169 was found by Koichi Itagati, Teppo-machi, Yamagata, Japan, on unfiltered CCD frames taken with a 0.60-m f/5.7 reflector (reported via S. Nakano, Sumoto, Japan) and by R. Arbour (cf. IAUC 8035). SN 2003ch in UGC 3787 and SN 2003ci in UGC 6212 were reported by B. Swift and W. Li (LOTOSS/KAIT; cf. IAUC 8094). The tabulated data for SN 2003cg are from Itagati, except the offset, which is computed from a measurement of the galaxy center by Y. Kushida (Yatsugatake South Base Observatory; CCD image taken by R. Kushida; communicated by Nakano).

SN  2003 UT  α_{2000}  δ_{2000}  Mag.  Offset
2003cg Mar. 21.51 10^{h}14^{m}15.^{s}97 + 3^{o}28′02′′.5 14.4 14′′ E, 5′′ N
2003ch Mar. 21.2 7 17 57.92 + 9 41 34.9 17.4 5′′.4 E, 13′′.4 N
2003ci Mar. 21.4 11 10 23.83 + 4 49 35.9 17.5 11′′ W, 11′′.9 S

Additional approximate unfiltered CCD magnitudes (from KAIT unless otherwise noted): SN 2003cg, Mar. 13.3, [19.0; 22.3, 14.9; 22.444, 14.3 (Itagati); 22.559, 15.5 (Kushida; through clouds); 22.835, 14.5 (Arbour). SN 2003ch, Mar. 10.2, [19.0; 22.2, 17.0. SN 2003ci, Mar. 5.4, [19.0; 22.4, 17.5. SN 2003cg was not present on earlier survey images taken by R. Kushida (her most recent taken on Mar. 8) and by Arbour (on Mar. 14), and Arbour adds that it is absent from the Digitized Sky Survey. Li provides position end figures 16^{s}.00, 1′′.7 for SN 2003cg. SN 1984E also appeared in NGC 3169.

IGR J16358–4726 M. Revnivtsev, Space Research Institute, Moscow, and Max-Planck-

Institut für Astrophysik, Garching; M. Tuerler, INTEGRAL Science Data Centre and Geneva Observatory; M. Del Santo, Space Astrophysics and Cosmic Physics Institute, Rome; N. J. Westergaard, Danish Space Research Institute, Copenhagen; N. Gehrels, Goddard Space Flight Center; and C. Winkler, Research and Science Support Department, ESTEC, Noordwijk, report the discovery with the INTEGRAL satellite (+ IBIS/ISGRI) on Mar. 19.4 UT of a variable source located at α = 16^{h}35^{m}.8, δ = −47^{o}26′ (equinox 2000.0; position uncertainty 1′.5). The source flux varied by a factor of ∼ 2 on time scales of hours. The average fluxes were ∼ 50 mCrab in the energy band 15–40 keV and 20 mCrab in the band 40–100 keV. Observations in other bands are encouraged.

2003 March 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green